# Know Me, Respond to Me: Benchmarking LLMs for Dynamic User Profiling and Personalized Responses at Scale

### v1

@@ 1 Introduction
[…]
@@ 3 Constructing Examples in [Uncaptioned image] PersonaMem At Scale
[…]
@@
We sample a set of random personas from PersonaHub (Ge et al., 2024), each comprising about one to three sentences, and augment them with additional demographic information and extended personal details. We also construct a timeline and populate it with events that align with the persona. These events serve as the general personal history, such as education, career development, and life experiences, to provide a richer context. The prompts used in the process can be found in Appendix C.
@@
To evaluate the quality of our generated data, we conduct a human study on 90 random query–response pairs from PersonaMem, each grounded in user persona, personal histories, and associated utterances in conversation. Three annotators assess each Q&A pair across four dimensions: appropriateness, relevance, correctness, and best response. Judgments were very high for all dimensions – 97.8%, 95.6%, 97.8%, and 90.0% respectively. Further details are provided in Appendix A.
@@ 4.2 Evaluating Language Models in Long-Context Settings
[…]
@@ 4.4 Evaluation with External Memory Modules
We evaluate whether using a retriever to identify relevant information in the history will help improve model’s performance. We evaluate two external memory approaches—RAG (Lewis et al., 2020) and Mem0 (Mem0, 2024)—against vanilla LLMs. For these experiments, we consider only the GPT-4o and GPT-4o-mini models.
@@ 4.5 Evaluation of Language Models in Generative Settings
[…]
@@ Results.
As shown in Figure 6, we observe the similar trend to our discriminative evaluation results in terms of difficulty by different user query types. Models get reasonably good performance on recalling facts and tracking preference changes, while giving new suggestions and generalizing to new scenarios are still the most challenging types of queries for models. Interestingly, when comparing the same model, specifically LLama-3.1-8B-instruct, under discriminative and generative settings, we see the performance is better in the generative setting, potentially suggesting that the model is able to provide a personalized response without seeing all the candidate options in the input. Since we only managed to run evaluation on 32k context length with the generative setting, it remains to be investigated whether results in generative vs. discriminative settings stand for longer context length and for different models. We also find that model performance declines as users’ new requests become more distant from their previously revealed information. Detailed results are provided in Appendix B.
@@ 5.2 Towards Personalization in Large Language Models
[…]
@@ 6 Conclusion
Appendix A Details on Human Evaluation
@@
We recruited three human annotators from within our institutions, who agreed to volunteer their time. All annotators are fluent English speakers working in academia. We iterated the annotation instructions and template with active feedback from the annotators, leading to the finalized version.
@@
Appendix B Supplementary Experiment Results
@@
Appendix C Prompts Used in [Uncaptioned image] PersonaMem  Dataset Generation
[…]
@@ Appendix C Prompts Used in [Uncaptioned image] PersonaMem  Dataset Generation
Generating task-specific user preferences. Here is the persona: {persona} Here are some events related to the person’s general background history: {general_personal_history} Given the persona above, please first list 20 hobbies related to {task}. Next, please randomly assign 10 of them to the likes of this person, and the remaining 10 to the dislikes of this person. Make sure every hobby, regardless of whether it is a like or dislike, is unique and attractive in common, so that the exact dislikes can potentially be turned into likes in the future. please list 10 unique personal hobbies and 10 things this person dislikes but others might still like, using bullet points, related to {task}. Next, write 10 more events related to the topic of {task}. Think about how this person’s general background history may affect their events under {task}. Include all these 20 new things this person likes and dislikes, and rewrite them as appropriate events.Do NOT mention anything already mentioned above. Do NOT mention anything about the general personal history, like the professional development. Each event must come with the related personal hobbies or dislikes, marked using a key ’[Fact] Likes:’ or ’[Fact] Dislikes:’ closely associated with the 20 things you listed here, and they should concentrate on the topic of{task}. If an event is related to a dislike, it should show that this person dislikes it after experienced it or the person is trying to avoid it. Use the same JSON format with MM/DD/YYYY timestamp from {start_time}, and use short-term/long-term labels as above. There should be 10 short-term and 10 long-term events.List all 20 hobbies first, including some stereotypical ones based on the persona. Mark stereotypical ones by square brackets ’[stereotypical]’. Next, randomly assign those 20 hobbies into likes or dislikes for this person. After you have generated the list above, generate one dict for each event following those 20 likes and dislikes. List all 20 hobbies first, and then follow this template in string to randomly assign those 20 hobbies into likes or dislikes for this person: 20 hobbies: xxx, …, xxx Initial preferences randomly assigned: [1] Likes xxx (Add [stereotypical] here if appropriate, same for each of the 20 rows below) [2] Likes xxx [3] Likes xxx [4] Likes xxx [5] Likes xxx [6] Likes xxx [7] Likes xxx [8] Likes xxx [9] Likes xxx [10] Likes xxx [1] Dislikes xxx [2] Dislikes xxx [3] Dislikes xxx [4] Dislikes xxx [5] Dislikes xxx [6] Dislikes xxx [7] Dislikes xxx [8] Dislikes xxx [9] Dislikes xxx [10] Dislikes xxx After you have generated the list above, here is the template in JSON you should follow for each event. PLEASE MUST USE JSON FOR THIS PART: ”MM/DD/YYYY”: ”Event”: xxx, ”Category”: ”Short-Term” OR ”Long-Term” ”[Fact] Likes” OR ”[Fact] Dislikes”: xxx, , Do NOT modify the names of these keys. Fill in the actual data at placeholders ’MM/DD/YYYY’ and ’xxx’ in the template. Please use DOUBLE quotes in order to generate the correct JSON format.

### v2

@@ 1 Introduction
PersonaMem features 20 personas, with over 180 interaction histories. Each interaction history contains 10, 20, or 60 sessions, where we dynamically adjust the total length of the history to approximately 32kk, 128kk, and 1MM tokens, respectively. Each session consists of 15–30 conversation turns between a user and an LLM chatbot. The user-LLM conversations span across 15 diverse topics, ranging from therapy and legal advice to recommendations on books, music, movies, and food; personal matters such as family, dating, health, and finance; and practical tasks like travel planning, online shopping, studying tips, and home decoration. In total, the benchmark features around 6kk in-situ user query and LLM response pairs across the 7 query types. Detailed dataset breakdown is discussed in Appendix D. The size of our benchmark is not limited by the scalability of the synthetic data pipeline but to make the evaluation cost reasonable.
@@ 3 Constructing Examples in [Uncaptioned image] PersonaMem At Scale
We develop a modular data curation pipeline—powered by GPT-4o (Hurst et al., 2024)—to synthesize persona-oriented, multi-session user–model conversations with long context. The pipeline minimizes irrelevant or randomly injected content to better evaluate how effectively LLM chatbots address the challenges outlined in Section 2, while ensuring cost-effectiveness and scalability: generating data for each persona on each conversation topic costs approximately \$2, independent of the context window length up to 1MM tokens.
@@
We sample a set of random personas from PersonaHub (Ge et al., 2024), each comprising about one to three sentences, and augment them with additional demographic information and extended personal details. We also construct a timeline and populate it with events that align with the persona. These events serve as the general personal history, such as education, career development, and life experiences, to provide a richer context. The prompts used in the process can be found in Appendix G.
@@
To evaluate the quality of our generated data, we conduct a human study on 90 random query–response pairs from PersonaMem, each grounded in user persona, personal histories, and associated utterances in conversation. Three annotators assess each Q&A pair across four dimensions: appropriateness, relevance, correctness, and best response. Judgments were very high for all dimensions – 97.8%, 95.6%, 97.8%, and 90.0% respectively. Further details are provided in Appendix B.
@@ 4.2 Evaluating Language Models in Long-Context Settings
We first evaluate language models in the long-context setting, where the full user-LLM interaction history is provided as input to the models. Due to the length of the history, all models here were evaluated zero-shot, without demonstration examples of other histories and user queries. Our evaluation covers GPT-4.1, o4-mini, o3-mini, GPT-4.5, o1, GPT-4o, GPT-4o-mini, Gemini-2.0-Flash, Gemini-2.0-Flash-Lite, Gemini-1.5-Flash, DeepSeek-R1-671B, Llama-4-Maverick, Llama-3.1-405B, Claude-3.7-Sonnet, and Claude-3.5-Haiku (OpenAI, 2025a; 2024b; b; 2024a; Hurst et al., 2024; Team et al., 2024; Guo et al., 2025; Grattafiori et al., 2024; Anthropic, 2024) on 128kk-token context windows. We also evaluate models that support longer contexts—Llama-4-Maverick, Gemini-2.0-Flash, Gemini-2.0-Flash-Lite, and Gemini-1.5-Flash—on 1MM-token context windows. We report the following findings:
@@ 4.4 Evaluation with External Memory Modules
We evaluate whether using a retriever to identify relevant information in the history will help improve model’s performance. We evaluate two external memory approaches—RAG (Lewis et al., 2020) and Mem0 (Mem0, 2024)—against vanilla LLMs. For these experiments, we consider only the GPT-4o and GPT-4o-mini models. We show their latency in Appendix E.
@@ 4.5 Evaluation of Language Models in Generative Settings
As the method requires logarithmic probability of output tokens over the entire vocabulary, which is often not available in proprietary models, we evaluate open-weight models—LLaMA-3.1–70B, LLaMA-3.1–8B, and DeepSeek-Distill-LLaMA–8B. Due to constraints in computation resources, we only evaluate the models on the 10-session version of the benchmark, which includes around 32kk-tokens per session.
@@ Results.
As shown in Figure 6, we observe the similar trend to our discriminative evaluation results in terms of difficulty by different user query types. Models get reasonably good performance on recalling facts and tracking preference changes, while giving new suggestions and generalizing to new scenarios are still the most challenging types of queries for models. Interestingly, when comparing the same model, specifically LLama-3.1-8B-instruct, under discriminative and generative settings, we see the performance is better in the generative setting, potentially suggesting that the model is able to provide a personalized response without seeing all the candidate options in the input. Since we only managed to run evaluation on 32k context length with the generative setting, it remains to be investigated whether results in generative vs. discriminative settings stand for longer context length and for different models. We also find that model performance declines as users’ new requests become more distant from their previously revealed information. Detailed results are provided in Appendix C.
@@ 5.2 Towards Personalization in Large Language Models
Turning to the dialogue setting, earlier works like LaMP and PersonaLLM consider personalization within a single turn or session (Salemi et al., 2023; Jiang et al., 2023; Kirk et al., 2024). More recently, ImplexConv (Li et al., 2025) focuses on modeling implicit reasoning within personalized conversations. PersonaBench (Tan et al., 2025) simulates social interactions among diverse users through numerous but shorter sessions and access to synthetic private user data. PersoBench (Afzoon et al., 2024) leverages existing persona-aware datasets to evaluate language quality, persona coverage, and consistency. LongLaMP (Kumar et al., 2024) focuses on generating long-form texts other than more interactive responses within long context. Zhao et al. (2025) introduce PrefEval, which evaluates LLMs’ preference-following abilities for 20 topics in persona-oriented dialogues of 100k+ tokens. PersonaMem, besides the flexible setting of generating numerous 1MM-token contexts efficiently, places greater emphasis on personas as simulated humans in user-model interactions, featuring multiple fine-grained personalization tasks where profiles and preferences evolve through temporally grounded events.
@@ 6 Conclusion
LOCOMO LongMemEval PrevEval PersonaMem
Focused Tasks Long-term memory Long-term memory User preferences Fine-grained personalized responses
Avg. Single Session Len 477 tokens 3k tokens No info 6k tokens
Max Context Len 9k tokens 1.5M tokens 100k tokens 1M tokens
Data Sources MSC & own ShareGPT & UltraChat & own LMSYS-Chat-1M PersonaHub & own
Query Perspective third-person first-person first-person first-person
Max # Knowledge Updates No update 1 No update 3
Multi-Session Reasoning Yes Yes No Yes
# LLMs Evaluated 4 5 6 15
Table 2: Comparison of related benchmarks, including LOCOMO (Maharana et al., 2024), LongMemEval (Wu et al., 2024), and PrefEval (Zhao et al., 2025). LOCOMO and LongMemEval focus on general long-term memory tasks. In contrast, PersonaMem centers on personalization beyond memory retrieval, with all conversations in our benchmark are built around a coherent user persona with evolving preferences, mimicking more realistic user-chatbot conversations. PrefEval, which focuses on personalization too, but by first generating user preferences and then inserting them into randomly sampled contexts.
@@
7 Acknowledgment
This work was supported by the National Science Foundation (NSF) under Grant CCF-2112665 (TILOS) and the Research Discretionary Fund from Camillo J. Taylor, and by the Office of Naval Research (ONR) MSU grant from Dan Roth.
Appendix A Limitations and future work
A.1 Broader context in user privacy concerns
Privacy is a critical aspect of LLM personalization in the real world. In our setting, we personalize responses based on only preferences and activities shared by the user in previous user-chatbot interactions, and the model uses this information for its own responses without external sharing. To avoid potential privacy risks associated with real user data, we intentionally propose a synthetic data curation pipeline in this work. This synthetic approach allows researchers in the community to safely explore personalization methods. One possible direction for future work could be designing question-answer pairs that specifically involve sensitive user information.
A.2 More advanced retrieval methods
Our current exploration of retrieval-augmented methods, such as RAG and Mem0, is intended as a proof of concept, as the primary focus of this work is on the design and release of the personalization benchmark. We are excited to encourage more exploration on state-of-the-art long-context, memory, and retrieval-augmented generation methods in future work, especially those that preserve and understand the evolution of user personas and reasons behind preference updates, as well as enhancing user personalization in new or unseen scenarios.
A.3 Potential artifacts in the synthetic data generation process
To reduce artifacts that might make the benchmark artificially easier, we’ve taken several steps. For example, we removed question-answer pairs where the correct answer was unintentionally obvious, such as being noticeably longer or sharing identical key words with the questions. We also filtered out queries that an LLM can answer correctly more than once in three attempts, without seeing any actual conversation context. Besides, we have included checks in our human evaluations to confirm that the correct answers can indeed be derived from the provided context.
A.4 Potential gaps between evaluations on open-ended generations and multiple choices
In purely open-ended generative settings, personalization can lead to many possible correct answers, depending on how the user persona is used and which related user preference is used. Meanwhile, open-ended evaluations are computationally expensive due to the need for LLM-as-a-Judge for each question-answer pair. As a result, we evaluate generative tasks by computing the joint log-likelihood of each candidate option, without explicitly presenting all four options in the prompt. This approach yields similar patterns with those observed in standard discriminative evaluations in our experiment, while offering a more reliable basis for benchmarking performance compared to fully open-ended ones.
Appendix B Details on Human Evaluation
@@
We recruited three authors from among the authors of this work. [^1] We recognize that the authors may not be fully impartial annotators. To reduce this issue, the three authors who participated were not directly involved with the data generation process. We nevertheless will consider external annotators for future work. We iterated the annotation instructions and template with active feedback from the annotators, leading to the finalized version.
@@
Appendix C Supplementary Experiment Results
@@
Appendix D Detailed Breakdown of the [Uncaptioned image] PersonaMem Statistics
Below is a more detailed breakdown of the dataset.
D.1 Different Query Types
  • Recall_user_shared_facts: 5.8%
  • Acknowledge_latest_user_preferences: 30.09%
  • Track_full_preference_evolution: 10.97%
  • Revisit_reasons_behind_preference_updates: 9.28%
  • Provide_preference_aligned_recommendations: 11.58%
  • Suggest_new_ideas: 22.92%
  • Generalize_to_new_scenarios: 9.35%
D.2 Different Conversation Topics
  • Book Recommendation: 6.3%
  • Dating Consultation: 7.2%
  • Family Relations: 5.3%
  • Financial Consultation: 7.3%
  • Food Recommendation: 8.4%
  • Home Decoration: 5.6%
  • Legal Consultation: 10.4%
  • Medical Consultation: 7.2%
  • Movie Recommendation: 5.8%
  • Music Recommendation: 1.6%
  • Online Shopping: 7.2%
  • Sports Recommendation: 7.2%
  • Study Consultation: 5.8%
  • Therapy: 9.1%
  • Travel Planning: 5.7%
Figure 11: Distribution of Query Types in the Dataset
Figure 12: Distribution of Conversation Topics in the Dataset
D.3 Distance from the User Query to the Reference Information in the Context (PersonaMem_128k)
  • 0-2 sessions: 5.6%
  • 3-6 sessions: 20.1%
  • 7-10 sessions: 17.6%
  • 11-14 sessions: 17.9%
  • 15-18 sessions: 23.6%
  • 19-20 sessions: 15.2%
D.4 Distance from the User Query to the Reference Information in the Context (PersonaMem_128k) in Tokens
  • 0-9.18k tokens: 5.7%
  • 9.18k-22.3k tokens: 14.8%
  • 22.3k-35.4k tokens: 11.3%
  • 35.4k-48.5k tokens: 7.4%
  • 48.5k-61.6k tokens: 8.2%
  • 61.6k-74.7k tokens: 8.1%
  • 74.7k-87.8k tokens: 8.6%
  • 87.8k-101k tokens: 11.6%
  • 101k-114k tokens: 17.1%
  • 114k-128k tokens: 7.3%
Figure 13: Session Distance from User Query to Reference Information
Figure 14: Token Distance from User Query to Reference Information
D.5 For PersonaMem_1M
D.5.1 Distance from the User Query to the Reference Information in the Context (PersonaMem_1M) in Terms of Sessions
  • 0-7 sessions: 5.6%
  • 8-13 sessions: 6.1%
  • 14-19 sessions: 10.1%
  • 20-25 sessions: 11.4%
  • 26-31 sessions: 8.3%
  • 32-37 sessions: 8.9%
  • 38-43 sessions: 9.6%
  • 44-49 sessions: 9.9%
  • 50-55 sessions: 11.7%
  • 56-60 sessions: 18.3%
D.5.2 Distance from the User Query to the Reference Information in the Context (PersonaMem_1M) in Tokens
  • 0-101k tokens: 6.1%
  • 101k-195k tokens: 5.5%
  • 195k-288k tokens: 10.3%
  • 288k-381k tokens: 10.2%
  • 381k-474k tokens: 12.8%
  • 474k-568k tokens: 8.3%
  • 568k-661k tokens: 9.1%
  • 661k-754k tokens: 9.6%
  • 754k-847k tokens: 11.4%
  • 847k-1M tokens: 16.7%
Figure 15: Session Distance from User Query to Reference Information
Figure 16: Token Distance from User Query to Reference Information
Appendix E The latency of the different approaches with external retrieval modules
In our experiment using GPT-4o-mini with a 32k-token context window and 589 user queries, RAG completed all queries in 6 minutes, averaging 0.61 seconds per query. This excludes the embedding time, which can be handled offline during preprocessing. RAG achieves constant-time retrieval, independent of context length. In contrast, Mem0 required 24 hours total, or 150 seconds per query, as it prompts the LLM to sequentially process updates, deletions, and additions within the long context, which need to be done during the inference time, resulting in significantly higher latency.
Appendix F Analysis of error patterns
We conducted a manual error analysis on 100 randomly selected user queries where GPT-4o failed to select the most personalized responses. We categorized the errors into the following five main types:
  • Format Error (14%) – The model fails to select a valid option from the provided choices.
  • Hallucination (12%) – The model selects an option that contains preferences never mentioned by the user.
  • Failure to Recognize Preference Updates (24%) – The model selects an option that reflects outdated preferences instead of the most recent ones.
  • Lack of Personalization (48%) – The model selects a generally reasonable option, instead of a more personalized one to the current user.
  • Other (2%) – Miscellaneous errors.
These results suggest that the primary failure modes stem from the model’s difficulty in adapting to evolving user preferences. Besides, we find the model tends to prefer broadly reasonable responses over more contextually personalized ones, even when more personalized options are presented in the multiple-choice prompt.
@@ Appendix F Analysis of error patterns
Appendix G Prompts Used in [Uncaptioned image] PersonaMem  Dataset Generation
Persona Description $\Rightarrow$ Initial General User Profile and Preferences Given the following persona, expand it with 10 person’s general background history within ten years starting at {start_time}. Turn each point into the format of a bullet point, and add a timestamp in the format of MM/DD/YYYY for each bullet point. Remember that these events should be general like career development, and they will be shared across multiple different topics.You should mention both daily activities and important key milestones, and both positive and negative history events. Also relate history to what this person prefers and dislikes. Use JSON format where each timestamp is a key in the JSON dictionary. Each point should also be marked with labels of either [’Short-Term’] or [’Long-Term’], where short-term fact refers to something happening daily, which can be irrelevant to the persona like what the person eats, which should come with temporal quantifiers like ’today’ or so, but long-term fact refers to some key personas that won’t be changed for at least a year. There should be 5 short-term and 5 long-term events. Include all 10 things this person likes and dislikes mentioned in the persona, and rewrite them as appropriate events. All events must have an appropriate time stamp in the format of MM/DD/YYYY. List at least 10 events, more are welcome. Here is the template you should follow for each event: ”MM/DD/YYYY”: { ”Event”: xxx, ”Category”: ”Short-Term” OR ”Long-Term” }, Do NOT modify the names of these keys. Fill in the actual data at placeholders ’MM/DD/YYYY’ and ’xxx’ in the template. Please use DOUBLE quotes in order to generate the correct JSON format. Here is the persona: {persona}
@@ Appendix G Prompts Used in [Uncaptioned image] PersonaMem  Dataset Generation
Generating task-specific user preferences. Here is the persona: {persona} Here are some events related to the person’s general background history: {general_personal_history}
@@
Given the persona above, please first list 20 hobbies related to {task}. Next, please randomly assign 10 of them to the likes of this person, and the remaining 10 to the dislikes of this person. Make sure every hobby, regardless of whether it is a like or dislike, is unique and attractive in common, so that the exact dislikes can potentially be turned into likes in the future. please list 10 unique personal hobbies and 10 things this person dislikes but others might still like, using bullet points, related to {task}. Next, write 10 more events related to the topic of {task}. Think about how this person’s general background history may affect their events under {task}.
Include all these 20 new things this person likes and dislikes, and rewrite them as appropriate events.Do NOT mention anything already mentioned above. Do NOT mention anything about the general personal history, like the professional development. Each event must come with the related personal hobbies or dislikes, marked using a key ’[Fact] Likes:’ or ’[Fact] Dislikes:’ closely associated with the 20 things you listed here, and they should concentrate on the topic of{task}. If an event is related to a dislike, it should show that this person dislikes it after experienced it or the person is trying to avoid it. Use the same JSON format with MM/DD/YYYY timestamp from {start_time}, and use short-term/long-term labels as above. There should be 10 short-term and 10 long-term events.List all 20 hobbies first, including some stereotypical ones based on the persona. Mark stereotypical ones by square brackets ’[stereotypical]’. Next, randomly assign those 20 hobbies into likes or dislikes for this person. After you have generated the list above, generate one dict for each event following those 20 likes and dislikes. List all 20 hobbies first, and then follow this template in string to randomly assign those 20 hobbies into likes or dislikes for this person: 20 hobbies: xxx, …, xxx Initial preferences randomly assigned: [1] Likes xxx (Add [stereotypical] here if appropriate, same for each of the 20 rows below) [2] Likes xxx [3] Likes xxx [4] Likes xxx [5] Likes xxx [6] Likes xxx [7] Likes xxx [8] Likes xxx [9] Likes xxx [10] Likes xxx [1] Dislikes xxx [2] Dislikes xxx [3] Dislikes xxx [4] Dislikes xxx [5] Dislikes xxx [6] Dislikes xxx [7] Dislikes xxx [8] Dislikes xxx [9] Dislikes xxx [10] Dislikes xxx After you have generated the list above, here is the template in JSON you should follow for each event. PLEASE MUST USE JSON FOR THIS PART: ”MM/DD/YYYY”: ”Event”: xxx, ”Category”: ”Short-Term” OR ”Long-Term” ”[Fact] Likes” OR ”[Fact] Dislikes”: xxx, , Do NOT modify the names of these keys. Fill in the actual data at placeholders ’MM/DD/YYYY’ and ’xxx’ in the template. Please use DOUBLE quotes in order to generate the correct JSON format.